# Modelling vehicle and pedestrian collective dynamics: Challenges and advancements

### v1

@@ 1. Introduction
Our societies are becoming increasingly urbanised. Transport and mobility are also fundamental to the quality of everyday life. In this context, managing traffic flow, mixed urban traffic, and pedestrian dynamics is of considerable interest for public safety, economic development, and the conservation of the environment. In particular, crowd management in large buildings, such as train stations or stadiums, or during large outdoor events, such as festivals or concerts, has become the focus of much attention [36]. However, controlling traffic flow and pedestrian dynamics is not straightforward. In fact, pedestrian crowding and traffic flow are collective behaviours of many-body systems subject to complex social and physical interactions. In practice, intelligent traffic systems and crowd management are generally designed based on simulations and computational optimisation techniques. Simulations are based on pedestrian and traffic flow modelling approaches, and the demand for reliable and realistic models is getting more and more pressing. However, much of the collective drivers’ and pedestrians’ behaviour remains poorly understood. For example, traffic flow stabilisation is still a challenge, especially in the context of driving automation [44], while the notion of anticipation for a pedestrian is still somewhat fuzzy.
In congested conditions, pedestrian and road traffic flows essentially describe collective dynamics and coordinated behaviours. These coordinated dynamics can result from a form of collective intelligence, for example, when pedestrian counterflows self-organise into lanes to optimise flow [47] (although this phenomenon is also observed in passive particle systems such as colloids with different directions of motion), when crowds clear the way to facilitate the passage of an intruder [93, 11], for example an ambulance (e.g., in the video [2]), or when pedestrians spread out over different possible exits to balance the load on each exit during evacuation [136]. That being said, collective dynamics are not always beneficial and can even be strongly detrimental. Consider the stop-and-go waves (accordion traffic) commonly observed on congested motorways, which negatively impact traffic safety and performance and heighten the environmental harm caused by traffic because of the increased gas consumption due to repeated accelerations and braking. For all these problems, modelling, understanding and controlling collective dynamics starting from the local interactions of the agents from a complex systems perspective is a challenging task [9].
@@ 1.1. Modelling vehicle and pedestrian collective dynamics
The modelling of vehicle and pedestrian collective dynamics is generally divided into three different levels: strategic, tactical, and operational. This concept, that originally finds its roots in the military field, was introduced to the modelling of driver behaviour in the 1970s by Michon [83] and was transferred to pedestrian dynamics in the early 2000s [54]. At the strategic level, the modeller is concerned with the largest scales, those of activity choices and scheduling (departure time planning). The focus of the tactical level is on smaller scales; it includes manoeuvre planning, the choice of a path, or the choice of an exit in the context of an evacuation. Operational modelling describes local motion in interaction with neighbours and obstacles. The state of the art generally treats the strategic, tactical and operational modelling scales separately. However, we argue in this review that these classifications are not entirely distinct from one another, but could be strongly interlaced, especially the operational and tactical modelling scales. This leads to conceptual problems in modelling these systems. Among operational models, three categories are classically distinguished: microscopic, mesoscopic, and macroscopic approaches [19, 46]. Microscopic models deal with the individual motion of the agents in Lagrangian coordinates, while mesoscopic and macroscopic approaches aggregate the dynamics in Eulerian coordinates using probabilistic agent distribution functions and averaged quantities such as flow and density, respectively. Among these approaches, first-order models posit that the agent’s speed instantaneously adapts to the environment, typically using proactive models based on anticipation and optimisation, whereas force-based models are second-order approaches that are formally similar to Newton’s equation and thus incorporate into the dynamics ‘inertia’, i.e., gradual adjustments of an agent’s velocity to the current situation, along with possible mechanical forces.
@@ 2. Vehicle and pedestrian models
We start by briefly reviewing continuous vehicle and pedestrian operational models using differential systems. The review includes car-following in one dimension in section 2.1 and pedestrian models in two dimensions in section 2.2, covering force-based reactive models and anticipation-based proactive approaches. We refer to the books [27, 79, 106, 127] for more in-depth reviews of the topic, including mesoscopic and macroscopic modelling approaches.
@@ 2.1.1. Pioneering follow-the-leader models (1950-1970)
The first studies on car-following behaviour can be traced back to the 1950s and the pioneering work by Reuschel on the dynamics of road vehicles [102]. In this first car-following model, the forerunner of a long series of works, the speed is directly proportional to distance to the vehicle in front
@@
where $\lambda$ is here a speed relaxation time. In this model, there is no regulation of distance as a function of velocity, and any uniform speed configuration is systematically stable. The end of the 1950s saw the fruitful development of delayed linear models and collective stability analysis [17, 51, 65]. Nonlinear models date back to early 1960s with, e.g., the general Stimulus-Response follow-the-leader model by Gazis et al. [41]
@@
where $\alpha$ and $\beta$ are some exponent parameters, and the first-order exponential delayed model of Newell [90]
@@
In fact, in congested traffic, the speed of the vehicles is assumed to be proportional to the spacing ahead, the time gap $T$ being the inverse of the proportionality factor. It corresponds to the time before collision between a vehicle and its predecessor if the latter were to stop abruptly while the speed of the follower remains constant. The constant-time-gap-policy is a following strategy that was initiated in the California Code in the US [17] and physically related to safety distances in the early 1980s [42]. The policy is now widely accepted for modelling driver behaviour and also for Adaptive Cruise Control (ACC) systems (see the pioneering work [116]) that is recommended in ISO standards (such as the ISO15622 with a time gap gap ranging from 0.8 to 2.2 s [55] or again 1.8 s for the German ADAC guidelines).
@@ 2.1.2. Modern car-following models (1990-today)
A second spate of nonlinear models emerged in the 1990s. In parallel, the terminology follow-the-leader was on the wane, in favour of car-following (see [24, Fig. 2]). This second spate is often referred to as Optimal Velocity (OV) models. The concept of Optimal Velocity comes from the systematic use of a nonlinear equilibrium speed function depending on the distance to the preceding vehicle, a notion initially introduced in the pioneering models by Reuschel [102], Chandler [17] and Newell [90]. Note that the terminology Optimal Velocity can be misleading, as there is no optimisation process here. That said, it has made its way into the literature. In modern OV models, the OV function $V:\mathbb{R}\mapsto\mathbb{R}_{+}$ can take the form of a sigmoidal arc-tangent as in the seminal Optimal Velocity model of Bando et al. [6]
@@
or the Full Velocity Difference (FVD) model by Jiang et al. [57]
@@
including two relaxation rates: $\tau_{1}$ is associated with the relaxation to the distance-dependent OV, whereas the speed difference with the predecessor relaxes over a characteristic time $\tau_{2}$. This double dependence is not surprising; it is sensible to assume that car-following behaviour depends not only on the distance to the preceding car, but also on their speed difference. Nowadays all modern car-following models depend on both of these variables and also on individual speeds, as inter alia in the famous Intelligent Driver (ID) model by Treiber et al.  [126]
@@
The emergence of new nonlinear models in the 1990s and early 2000s notably followed from simulation results showing realistic stop-and-go waves in stationary states for fine-tuned parameter settings [6, 5, 57]. The mathematical analysis of the models yielded collective stability conditions, typically $2\tau<1/V^{\prime}(d)=:T$ for the OV model [6]. In other words, the reaction time $\tau$ must be sufficiently small compared to the time gap $T$, that is, the drivers must be sufficiently reactive. Such findings opened the door to the modelling and control of traffic jam and stop-and-go dynamics [139, 96]. A more recent revival comes from the advent of driving automation and the need to develop stable models [22, 44]. For example, the Adaptive Time Gap model [122], where the time gap $T_{n}(t)=\big{[}\Delta x_{n}(t)-\ell\big{]}/\dot{x}_{n}(t)$, which is the distance gap divided by the velocity, is relaxed to the desired time gap parameter $T$, i.e, $\dot{T}_{n}(t)=\frac{1}{\tau}[T-T_{n}(t)]$, or again in Newtonian notation
@@ 2.2. Pedestrian models
Turning to pedestrians, the world gains a new geometric dimension, as pedestrians do not evolve on predefined tracks. Thus, pedestrian models are two-dimensional models that describe the dynamics of the pedestrian position $\vec{x}_{n}(t)\in\mathbb{R}^{2}$ according to the positions and velocities of its neighbours. In contrast to car-following models which have nearest neighbour interaction schemes in one dimension, the modelled pedestrians typically interact with all their close neighbours in two dimensions. The interactions may depend on distance, velocity difference, and also safety related variables such as the time-to-collision (TTC). Pioneering models are based on a superposition of pairwise interaction forces with the neighbours, the so-called social forces [50], while more recent approaches rely on collision avoidance, anticipation, and optimisation concepts. Pedestrians are generally represented as disks, as in Figure 2, although more realistic descriptions use ellipses or even skeleton models for the pedestrian shapes.
$\ell$$\vec{e}_{n,p}$$\dot{\vec{x}}_{n}$$\vec{x}_{n}$$\vec{x}_{m}$$\vec{x}_{p}$$|\Delta\vec{x}_{n,m}|$
@@ 2.2.1. Force-based models
The concept of a social force field can be traced back to the seminal works of Lewin in the social psychology of group dynamics [68, 69, 20]. Force-based pedestrian models date back to the pioneering work of Hirai and Tarui in the 1970s on the simulation of crowd behaviour in panic [52] where the dynamics of a pedestrian result from the superposition of different pairwise forces, including attraction and repulsion forces with the neighbours and velocity alignment force. The concept of social force for pedestrian dynamics was later democratised from the 1990s with the work of Helbing and co-authors [50, 48]. Around the same time, swarming-type models such as the Boids model by Reynolds [103] and self-propelled particle model by Vicsek and co-authors [134], or, later, Cucker-Smale and other swarming and flock models [30, 15] have been developed with interaction mechanisms based on averages with neighbours and similar principles of pairwise interaction superposition. However, in contrast to swarming models, space is not isotropic in pedestrian models: agents experience a drift towards their desired destinations in addition to the interaction field.
@@
In the Centrifugal Force (CF) model [144] and the Generalised Centrifugal Force (GCF) model [21], the range of interaction decays algebraically. The repulsive force for the CF model is
@@
where $\beta,\eta>0$ are parameters, $r\in C(\mathbb{R}^{2},R_{+})$ models a velocity-dependent pedestrian shape, e.g. an ellipse [21], while $\tilde{\Delta}\dot{\vec{x}}$ is the relative velocity projected on the direction to the neighbouring pedestrians. Unlike the SF model, the CF and GCF models also rely on the relative velocity to the neighbours. Nowadays, many extensions of the SF model include relative velocity terms and other velocity-dependent mechanisms; see the review [18] and references therein. For instance, in [98], the SF model extension consists of some respect areas in the vicinity of the pedestrians that force them to stop. Similarly, but in a dynamic way, in [145] the SF model is extended with explicit collision prediction. In another recent approach, the interaction force is anisotropic to take into account for a preferred crossing direction [125].
@@ 2.2.2. Velocity-based models
For pedestrian models, first-order models came second to second-order ones: A stream of models developed subsequently to the force-based models are the velocity-based models. In contrast to their Newton-like forebears, velocity-based models instantly regulate velocity at the first order. They are inherently free of inertia and allow strict control of body exclusion - an aspect that is difficult to manage with a force-based model [20, 113]. Velocity-based models are initially borrowed from the robotics. They generally rely on velocity optimisation processes under collision constraints over given anticipation times.
@@
Such developments can be traced back to the pioneering work of Fiorini et al. using purely geometric Velocity Obstacle (VO) models [37]. The model is based on linear extrapolations of the pedestrian trajectories to determine potential collision sets. Geometrically, these sets assume the form of so called collision cones under the hypothesis of constant pedestrian velocities. The cone is empty if no collision is expected. The velocity dynamics of agent $n$ is then obtained by minimising the deviation from the desired velocity under the exclusion constraint of the time-dependent collision cones $\mathcal{C}_{n}^{m}(t)$ with other agents $m$
(15) $\dot{\vec{x}}_{n}(t)=\text{arg}\min_{\vec{v}_{n}\,\not\in\,\cup_{m\neq n}\mathcal{C}_{n}^{m}(t)}\|\vec{v}_{0}-\vec{v}_{n}\|^{2}$
@@
where $\vec{v}_{0}$ is the desired velocity. Since the collision cone depends on the current velocities of the agents, the velocity model is implicit. In practice, the implicit system is solved numerically using semi-implicit schemes combining an explicit Euler solver for the velocity with an implicit Euler scheme for the positions of the agents. Initially, velocity obstacle models were developed for controlling robots in an environment including moving obstacles and automated vehicles driving on a highway [37]. First applications for pedestrian dynamics date back to the late 2000s [97, 131].
@@
Generally speaking, velocity obstacle models can be formulated as an optimisation problem under constraints [133]
(16) $\dot{\vec{x}}_{n}(t)=\text{arg}\min_{\vec{v}_{n}\,\not\in\,\cup_{m\neq n}\mathcal{C}_{n}^{m}(t)}C_{\vec{x}(t)}^{\vec{v}_{0}}(\vec{v}_{n})$
@@
where $C_{\vec{x}}^{\vec{v}_{0}}$ is a cost function depending on the desired velocity and also the positions and velocities of the neighbours. In the collision-free model developed in [97], the cost function includes components associated with speed variations, distance to desired speed, and orientation changes. In the Reciprocal Velocity Obstacle (RVO) model [131], the cost is the sum of a quadratic deviation from the desired velocity as in the original velocity obstacle model plus a repulsive potential based on the time to collision. The cost function is a quadratic deviation from the desired speed as in the initial VO model (15) in the Optimal Reciprocal Collision Avoidance (ORCA) model [130], while its extension PORCA model [75] includes an additional velocity-dependent term that prevents the system from freezing. No collision cones are used in the Linear Trajectory Avoidance model [99]. Instead, the optimisation is done by exponentially penalising the minimum distance under linear extrapolation to avoid collisions, in addition to a quadratic deviation from the desired speed as in the original VO model (15).
@@
A general collision-free velocity-based modelling framework was proposed in the 2000s by Maury and Venel [80, 81]. Its principle also rests on velocity optimisation under collision-free constraints. However, unlike the models described above where the pedestrian velocities are optimised individually, the optimisation is performed centrally over all the pedestrians. The pedestrians are considered as disks of diameter $\ell$. The set of admissible configurations of the system excluding overlapping configurations is given by

### v2

# Modelling vehicle and pedestrian collective dynamics: Challenges and advances

@@ 1. Introduction
Our societies are becoming increasingly urbanised. Transport and mobility are also fundamental to the quality of everyday life. In this context, managing traffic flow, mixed urban traffic, and pedestrian dynamics is of considerable interest for public safety, economic development, and the conservation of the environment. In particular, crowd management in large buildings, such as train stations or stadiums, or during large outdoor events, such as festivals or concerts, has become the focus of much attention [feliciani2022introduction]. However, controlling traffic flow and pedestrian dynamics is not straightforward. In fact, pedestrian crowding and traffic flow are collective behaviours of many-body systems subject to complex social and physical interactions. In practice, intelligent traffic systems and crowd management are generally designed based on simulations and computational optimisation techniques. Simulations are based on pedestrian and traffic flow modelling approaches, and the demand for reliable and realistic models is getting more and more pressing. However, much of the collective drivers’ and pedestrians’ behaviour remains poorly understood. For example, traffic flow stabilisation is still a challenge, especially in the context of driving automation [gunter2020commercially], while the notion of anticipation for a pedestrian is still somewhat fuzzy.
In congested conditions, pedestrian and road traffic flows essentially describe collective dynamics and coordinated behaviours. These coordinated dynamics can result from a form of collective intelligence, for example, when pedestrian counterflows self-organise into lanes to optimise flow [helbing2005self] (although this phenomenon is also observed in passive particle systems such as colloids with different directions of motion), when crowds clear the way to facilitate the passage of an intruder [nicolas2019mechanical, bonnemain2023pedestrians], for example an ambulance (e.g., in the video [MadrasAmbulance]), or when pedestrians spread out over different possible exits to balance the load on each exit during evacuation [wagoum2017understanding]. That being said, collective dynamics are not always beneficial and can even be strongly detrimental. Consider the stop-and-go waves (accordion traffic) commonly observed on congested motorways, which negatively impact traffic safety and performance and heighten the environmental harm caused by traffic because of the increased gas consumption due to repeated accelerations and braking. For all these problems, modelling, understanding and controlling collective dynamics starting from the local interactions of the agents from a complex systems perspective is a challenging task [bellomo2012modeling].
@@ 1.1. Modelling vehicle and pedestrian collective dynamics
The modelling of vehicle and pedestrian collective dynamics is generally divided into three different levels: strategic, tactical, and operational. This concept, that originally finds its roots in the military field, was introduced to the modelling of driver behaviour in the 1970s by Michon [michon1985critical] and was transferred to pedestrian dynamics in the early 2000s [hoogendoorn2004pedestrian]. At the strategic level, the modeller is concerned with the largest scales, those of activity choices and scheduling (departure time planning). The focus of the tactical level is on smaller scales; it includes manoeuvre planning, the choice of a path, or the choice of an exit in the context of an evacuation. Operational modelling describes local motion in interaction with neighbours and obstacles. The state of the art generally treats the strategic, tactical and operational modelling scales separately. However, we argue in this review that these classifications are not entirely distinct from one another, but could be strongly interlaced, especially the operational and tactical modelling scales. This leads to conceptual problems in modelling these systems. Among operational models, three categories are classically distinguished: microscopic, mesoscopic, and macroscopic approaches [chowdhury2000statistical, helbing2001traffic]. Microscopic models deal with the individual motion of the agents in Lagrangian coordinates, while mesoscopic and macroscopic approaches aggregate the dynamics in Eulerian coordinates using probabilistic agent distribution functions and averaged quantities such as flow and density, respectively. Among these approaches, first-order models posit that the agent’s speed instantaneously adapts to the environment, typically using proactive models based on anticipation and optimisation, whereas force-based models are second-order approaches that are formally similar to Newton’s equation and thus incorporate into the dynamics ‘inertia’, i.e., gradual adjustments of an agent’s velocity to the current situation, along with possible mechanical forces.
@@ 2. Vehicle and pedestrian models
We start by briefly reviewing continuous vehicle and pedestrian operational models using differential systems. The review includes car-following in one dimension in section 2.1 and pedestrian models in two dimensions in section 2.2, covering force-based reactive models and anticipation-based proactive approaches. We refer to the books [cristiani2014multiscale, maury2018crowds, schadschneider2010stochastic, treiber2013traffic] for more in-depth reviews of the topic, including mesoscopic and macroscopic modelling approaches.
@@ 2.1.1. Pioneering follow-the-leader models (1950-1970)
The first studies on car-following behaviour can be traced back to the 1950s and the pioneering work by Reuschel on the dynamics of road vehicles [reuschel1950fahrzeugbewegungen]. In this first car-following model, the forerunner of a long series of works, the speed is directly proportional to distance to the vehicle in front
@@
where $\lambda$ is here a speed relaxation time. In this model, there is no regulation of distance as a function of velocity, and any uniform speed configuration is systematically stable. The end of the 1950s saw the fruitful development of delayed linear models and collective stability analysis [chandler1958traffic, herman1959traffic, kometani1958stability]. Nonlinear models date back to early 1960s with, e.g., the general Stimulus-Response follow-the-leader model by Gazis et al. [gazis1961nonlinear]
@@
where $\alpha$ and $\beta$ are some exponent parameters, and the first-order exponential delayed model of Newell [newell1961nonlinear]
@@
In fact, in congested traffic, the speed of the vehicles is assumed to be proportional to the spacing ahead, the time gap $T$ being the inverse of the proportionality factor. It corresponds to the time before collision between a vehicle and its predecessor if the latter were to stop abruptly while the speed of the follower remains constant. The constant-time-gap-policy is a following strategy that was initiated in the California Code in the US [chandler1958traffic] and physically related to safety distances in the early 1980s [gipps1981behavioural]. The policy is now widely accepted for modelling driver behaviour and also for Adaptive Cruise Control (ACC) systems (see the pioneering work [swaroop1994comparision]) that is recommended in ISO standards (such as the ISO15622 with a time gap gap ranging from 0.8 to 2.2 s [iso2018intelligent] or again 1.8 s for the German ADAC guidelines).
@@ 2.1.2. Modern car-following models (1990-today)
A second spate of nonlinear models emerged in the 1990s. In parallel, the terminology follow-the-leader was on the wane, in favour of car-following (see [cordes2023single, Fig. 2]). This second spate is often referred to as Optimal Velocity (OV) models. The concept of Optimal Velocity comes from the systematic use of a nonlinear equilibrium speed function depending on the distance to the preceding vehicle, a notion initially introduced in the pioneering models by Reuschel [reuschel1950fahrzeugbewegungen], Chandler [chandler1958traffic] and Newell [newell1961nonlinear]. Note that the terminology Optimal Velocity can be misleading, as there is no optimisation process here. That said, it has made its way into the literature. In modern OV models, the OV function $V:\mathbb{R}\mapsto\mathbb{R}_{+}$ can take the form of a sigmoidal arc-tangent as in the seminal Optimal Velocity model of Bando et al. [bando1995dynamical]
@@
or the Full Velocity Difference (FVD) model by Jiang et al. [jiang2001full]
@@
including two relaxation rates: $\tau_{1}$ is associated with the relaxation to the distance-dependent OV, whereas the speed difference with the predecessor relaxes over a characteristic time $\tau_{2}$. This double dependence is not surprising; it is sensible to assume that car-following behaviour depends not only on the distance to the preceding car, but also on their speed difference. Nowadays all modern car-following models depend on both of these variables and also on individual speeds, as inter alia in the famous Intelligent Driver (ID) model by Treiber et al. [treiber2000congested]
@@
The emergence of new nonlinear models in the 1990s and early 2000s notably followed from simulation results showing realistic stop-and-go waves in stationary states for fine-tuned parameter settings [bando1995dynamical, bando1998analysis, jiang2001full]. The mathematical analysis of the models yielded collective stability conditions, typically $2\tau<1/V^{\prime}(d)=:T$ for the OV model [bando1995dynamical]. In other words, the reaction time $\tau$ must be sufficiently small compared to the time gap $T$, that is, the drivers must be sufficiently reactive. Such findings opened the door to the modelling and control of traffic jam and stop-and-go dynamics [wilson2011car, orosz2010traffic]. A more recent revival comes from the advent of driving automation and the need to develop stable models [ciuffo2021requiem, gunter2020commercially]. For example, the Adaptive Time Gap model [tordeux2010adaptive], where the time gap $T_{n}(t)=\big[\Delta x_{n}(t)-\ell\big]/\dot{x}_{n}(t)$, which is the distance gap divided by the velocity, is relaxed to the desired time gap parameter $T$, i.e, $\dot{T}_{n}(t)=\frac{1}{\tau}[T-T_{n}(t)]$, or again in Newtonian notation
@@ 2.2. Pedestrian models
Turning to pedestrians, the world gains a new geometric dimension, as pedestrians do not evolve on predefined tracks. Thus, pedestrian models are two-dimensional models that describe the dynamics of the pedestrian position $\vec{x}_{n}(t)\in\mathbb{R}^{2}$ according to the positions and velocities of its neighbours. In contrast to car-following models which have nearest neighbour interaction schemes in one dimension, the modelled pedestrians typically interact with all their close neighbours in two dimensions. The interactions may depend on distance, velocity difference, and also safety related variables such as the time-to-collision (TTC). Pioneering models are based on a superposition of pairwise interaction forces with the neighbours, the so-called social forces [helbing1995social], while more recent approaches rely on collision avoidance, anticipation, and optimisation concepts. Pedestrians are generally represented as disks, as in Figure 2, although more realistic descriptions use ellipses or even skeleton models for the pedestrian shapes.
[…]
@@ 2.2.1. Force-based models
The concept of a social force field can be traced back to the seminal works of Lewin in the social psychology of group dynamics [lewin1952field, lewin1968conceptual, chraibi2011force]. Force-based pedestrian models date back to the pioneering work of Hirai and Tarui in the 1970s on the simulation of crowd behaviour in panic [hirai1975simulation] where the dynamics of a pedestrian result from the superposition of different pairwise forces, including attraction and repulsion forces with the neighbours and velocity alignment force. The concept of social force for pedestrian dynamics was later democratised from the 1990s with the work of Helbing and co-authors [helbing1995social, helbing2000simulating]. Around the same time, swarming-type models such as the Boids model by Reynolds [reynolds1987flocks] and self-propelled particle model by Vicsek and co-authors [vicsek1995novel], or, later, Cucker-Smale and other swarming and flock models [cucker2007emergent, carrillo2010particle] have been developed with interaction mechanisms based on averages with neighbours and similar principles of pairwise interaction superposition. However, in contrast to swarming models, space is not isotropic in pedestrian models: agents experience a drift towards their desired destinations in addition to the interaction field.
@@
In the Centrifugal Force (CF) model [yu2005centrifugal] and the Generalised Centrifugal Force (GCF) model [chraibi2010generalized], the range of interaction decays algebraically. The repulsive force for the CF model is
@@
where $\beta,\eta>0$ are parameters, $r\in C(\mathbb{R}^{2},R_{+})$ models a velocity-dependent pedestrian shape, e.g. an ellipse [chraibi2010generalized], while $\tilde{\Delta}\dot{\vec{x}}$ is the relative velocity projected on the direction to the neighbouring pedestrians. Unlike the SF model, the CF and GCF models also rely on the relative velocity to the neighbours. Nowadays, many extensions of the SF model include relative velocity terms and other velocity-dependent mechanisms; see the review [chen2018social] and references therein. For instance, in [parisi2009modification], the SF model extension consists of some respect areas in the vicinity of the pedestrians that force them to stop. Similarly, but in a dynamic way, in [zanlungo2011social] the SF model is extended with explicit collision prediction. In another recent approach, the interaction force is anisotropic to take into account for a preferred crossing direction [totzeck2020anisotropic].
@@ 2.2.2. Velocity-based models
For pedestrian models, first-order models came second to second-order ones: A stream of models developed subsequently to the force-based models are the velocity-based models. In contrast to their Newton-like forebears, velocity-based models instantly regulate velocity at the first order. They are inherently free of inertia and allow strict control of body exclusion - an aspect that is difficult to manage with a force-based model [chraibi2011force, sticco2020effects]. Velocity-based models are initially borrowed from the robotics. They generally rely on velocity optimisation processes under collision constraints over given anticipation times.
@@
Such developments can be traced back to the pioneering work of Fiorini et al. using purely geometric Velocity Obstacle (VO) models [fiorini1998motion]. The model is based on linear extrapolations of the pedestrian trajectories to determine potential collision sets. Geometrically, these sets assume the form of so called collision cones under the hypothesis of constant pedestrian velocities. The cone is empty if no collision is expected. The velocity dynamics of agent $n$ is then obtained by minimising the deviation from the desired velocity under the exclusion constraint of the time-dependent collision cones $\mathcal{C}_{n}^{m}(t)$ with other agents $m$
(15) $\dot{\vec{x}}_{n}(t)=\text{arg}\hskip-14.22636pt\min_{\vec{v}_{n}\,\not\in\,\cup_{m\neq n}\mathcal{C}_{n}^{m}(t)}\|\vec{v}_{0}-\vec{v}_{n}\|^{2}$
@@
where $\vec{v}_{0}$ is the desired velocity. Since the collision cone depends on the current velocities of the agents, the velocity model is implicit. In practice, the implicit system is solved numerically using semi-implicit schemes combining an explicit Euler solver for the velocity with an implicit Euler scheme for the positions of the agents. Initially, velocity obstacle models were developed for controlling robots in an environment including moving obstacles and automated vehicles driving on a highway [fiorini1998motion]. First applications for pedestrian dynamics date back to the late 2000s [paris2007pedestrian, van2008reciprocal].
@@
Generally speaking, velocity obstacle models can be formulated as an optimisation problem under constraints [van2020generalized]
(16) $\dot{\vec{x}}_{n}(t)=\text{arg}\hskip-14.22636pt\min_{\vec{v}_{n}\,\not\in\,\cup_{m\neq n}\mathcal{C}_{n}^{m}(t)}C_{\vec{x}(t)}^{\vec{v}_{0}}(\vec{v}_{n})$
@@
where $C_{\vec{x}}^{\vec{v}_{0}}$ is a cost function depending on the desired velocity and also the positions and velocities of the neighbours. In the collision-free model developed in [paris2007pedestrian], the cost function includes components associated with speed variations, distance to desired speed, and orientation changes. In the Reciprocal Velocity Obstacle (RVO) model [van2008reciprocal], the cost is the sum of a quadratic deviation from the desired velocity as in the original velocity obstacle model plus a repulsive potential based on the time to collision. The cost function is a quadratic deviation from the desired speed as in the initial VO model (15) in the Optimal Reciprocal Collision Avoidance (ORCA) model [van2011reciprocal], while its extension PORCA model [luo2018porca] includes an additional velocity-dependent term that prevents the system from freezing. No collision cones are used in the Linear Trajectory Avoidance model [pellegrini2009you]. Instead, the optimisation is done by exponentially penalising the minimum distance under linear extrapolation to avoid collisions, in addition to a quadratic deviation from the desired speed as in the original VO model (15).
@@
A general collision-free velocity-based modelling framework was proposed in the 2000s by Maury and Venel [maury2008mathematical, maury2011discrete]. Its principle also rests on velocity optimisation under collision-free constraints. However, unlike the models described above where the pedestrian velocities are optimised individually, the optimisation is performed centrally over all the pedestrians. The pedestrians are considered as disks of diameter $\ell$. The set of admissible configurations of the system excluding overlapping configurations is given by